# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

### v1

@@ 6.1 Low-frequency noise and breaking waves
The lower frequency signal and noise of the acceleration are amplified compared to the higher frequencies through the integration, Eqs. (2) and (3). Displacement measurements from IMUs are therefore inherently sensitive to low-frequency noise. The low-frequency noise can be seen in Figure 2 and 9, and is similar to spectra measured by other IMU-based buoys (Nose et al., 2023; Chuang et al., 2013; Arraigada and Partl, 2006; Collins et al., 2014; Brown et al., 2018; Veras Guimarães et al., 2018; Rabault et al., 2020; Rainville, 2022). The longest wave periods that can be measured robustly using IMUs are about $20$ seconds ($0.05$ Hz), depending on the conditions (Brown et al., 2018; Van Essen et al., 2018; Datawell BV, 2023). Note that up to $100$ s periods can be observed by using e.g. GPS to measure wave orbital velocities, although this is also not trivial (Thomson et al., 2018). Note that when comparing spectra the length of the signal should be similar in order to expect similar signal to noise ratio.
Consequently, the acceleration must be high-pass filtered with a chosen cut-off frequency. Choosing this cut-off frequency automatically is not trivial, since the assumed low-frequency noise often extends higher than $0.05$ Hz and can be difficult to distinguish from the signal. In this analysis we use a relatively simple algorithm where we attempt to find a minimum in the smoothed spectrum after $0.05$ Hz (Nose et al., 2018; Rabault et al., 2022) which is used as the cut-off frequency (see Sec. 33.1).
@@
We observe that the low-frequency noise occurs more frequently in the surf, and only sporadically in the open ocean. As wave breaking is more violent and happens much more often in the surf zone, it seems likely that high-frequency acceleration impulses from breaking waves are the cause of the low-frequency noise in the measured elevation. Specifically, rapid acceleration and rotation caused by a breaking wave insufficiently sampled into the fusion algorithm would cause the buoy to lose orientation and would no longer correctly extract the vertical acceleration. To test this hypothesis we configured buoys with different IMU sample rates ($208$ Hz and $52$ Hz), but equal output rate (filtered using a FIR-filter and decimated). In the wave flume (Sec. 3), both the low-frequency and the high-frequency buoys performed well on harmonic waves (Figure 3). However, when the buoys are subjected to breaking waves in the wave flume (Sec. 33.3), the buoy with the higher IMU-sampling (B25) performs far better than the other (B23), with low-frequency oscillations clearly visible.
@@
Additionally, a buoy in breaking waves requires much higher sample rates and ranges in measuring rotation and acceleration. Given the large accelerations measured in breaking waves (several $g$ in the open water, and more than $10g$ in the surf in our case (Sec. 5), and 8148--148 - - 14 reported by Sinclair (2014), Brown et al. (2018), and Feddersen et al. (2023a)), a much higher sample rate is probably required to have a chance at measuring these signals with any success.
In Sec. 5, Fig. 15(d) we were able to retrieve the elevation with a cut-off frequency as low as $0.05$ Hz.
@@ 6.2 Capturing and quantifying breaking waves
Feddersen et al. (2023a) captures the trajectory of a breaking wave generated at the Surf Ranch remarkably well. In this work we attempt to capture the trajectory in the waves breaking onto a beach (Sec. 5). The waves breaking on the beach do not consist of a single reproducible, relatively clean, wave. Yet, we do capture the trajectory well compared to the spectrum outside the surf and the photographs. The signal was filtered with a cut-off frequency as low as $0.05$ Hz, without resulting in exaggerated displacements. To our knowledge a breaking wave in the ocean has not been captured with a buoy in this way before.
@@ 7 Conclusions
The SFY, a small, light-weight, drifting or moored wave buoy design with the capability to detect breaking events has been described. The buoy is designed to work in coastal environments and communicates over the cellular network as well as stores the data on SD-card. Its design, code, and hardware is open source and available publicly [^4] https://github.com/gauteh/sfy. The buoy has been proven to work reliably up to $80$ km from the nearest cellular tower, and in its small form measures frequencies from $0.05$ Hz to about $2.2$ Hz (or $1.1$ Hz for the moored buoy, without correction). The main advantages of this buoy are:
@@
  • high-bandwidth: with a greater bandwidth for data transmission, the full time series of AHRS-processed high-frequency accelerations ($52$ Hz) is transmitted back. This allows a much more detailed on-land analysis compared to transmitting only the wave spectra.
  • open source: anyone can build and modify the buoy to adapt it to their experiment.
@@
The most challenging measurements were performed in the surf zone. The irregular waves breaking on the beach were not produced by a singular wave component or soliton. Yet, by measuring at very high frequency ($208$ Hz) and with a large dynamic range (±$\pm 16g$ and ±$\pm 1000$ dps), we were able to reconstruct the trajectories of the buoy through breaking and non-breaking waves on the beach with a higher degree of success than previously done. Necessary to this is that we were able to use a low cut-off frequency of $0.05$ Hz here, normally reserved for open-waters. The lack of exaggerated elevation amplitudes after integrating to displacement gives us confidence that the buoy measures correctly. In Table 4 we recommend parameter ranges depending on the deployment scenario.
@@
An important conclusion which we surmise will be valid for all IMU-based buoys is that low-frequency noise can to a large degree be explained by impulsive events, like breakers, that are not adequately sampled. The inadequate sampling causes the AHRS algorithm to become confused, and the large, but very short, acceleration peaks are not well captured. We find that by sampling the IMU and running the AHRS-algorithm at a higher frequency before downsampling its output (if bandwidth is limited), much better noise characteristics are achieved. By this logic, better noise characteristics or estimates based on the spectrum (like $H_{m0}$) may improve if impulsive events can be detected and excluded from the spectrum.
The advent of small, disposable drifting wave buoys with telemetry opens new opportunities for dense sampling of the surf zone and regions with strong spatial variability, such as tidal currents in open-ocean conditions (see, e.g., the studies of the wave modulation by the Moskenes tidal current in Northern Norway by Saetra et al. 2021; Halsne et al. 2022, 2023). Deploying large, densely spaced, arrays of buoys that sample at rates high enough to identify breakers and the accelerations in breaking events should allow for experiments that are currently not possible, or prohibitively impractical and expensive to do.

### v2

@@ 6.1 Low-frequency noise and breaking waves
The lower frequency signal and noise of the acceleration are amplified compared to the higher frequencies through the integration, Eqs. (2) and (3). Displacement measurements from IMUs are therefore inherently sensitive to low-frequency noise. The low-frequency noise can be seen in Figs. 2 and 9, and is similar to spectra measured by other IMU-based buoys (Nose et al., 2023; Chuang et al., 2013; Arraigada and Partl, 2006; Collins et al., 2014; Brown et al., 2018; Veras Guimarães et al., 2018; Rabault et al., 2020; Rainville, 2022). The longest wave periods that can be measured robustly using IMUs are about $20$ seconds ($0.05$ Hz), depending on the conditions (Brown et al., 2018; Van Essen et al., 2018; Datawell BV, 2023). Note that up to $100$ s periods can be observed by using e.g. GPS to measure wave orbital velocities, although this is also not trivial (Thomson et al., 2018; Collins et al., 2024). Note that when comparing spectra the length of the signal should be similar in order to expect similar signal to noise ratio.
[…]
@@
We observe that the low-frequency noise occurs more frequently in the surf, and only sporadically in the open ocean. As wave breaking is more violent and happens much more often in the surf zone, it seems likely that high-frequency acceleration impulses from breaking waves are the cause of the low-frequency noise in the measured elevation. Specifically, rapid acceleration and rotation caused by a breaking wave insufficiently sampled into the fusion algorithm would cause the buoy to lose orientation and would no longer correctly extract the vertical acceleration. To test this hypothesis we configured buoys with different IMU sample rates ($208$ Hz and $52$ Hz), but equal output rate (filtered using a FIR-filter and decimated). In the wave flume (Sec. 3), both the low-frequency and the high-frequency buoys performed well on harmonic waves (Fig. 3). However, when the buoys are subjected to breaking waves in the wave flume (Sec. 33.3), the buoy with the higher IMU-sampling (B25) performs far better than the other (B23), with low-frequency oscillations clearly visible.
In fact, a similar phenomenon is observed in GPS-based buoys where discontinuities in the time-series manifest as low-frequency noise (Björkqvist et al., 2016; Collins et al., 2024).
@@
[…]
In Sec. 5, Fig. 16d we were able to retrieve the elevation with a cut-off frequency as low as $0.04$ Hz.
@@ 6.2 Capturing and quantifying breaking waves
Feddersen et al. (2023a) captures the trajectory of a breaking wave generated at the Surf Ranch remarkably well. In this work we attempt to capture the trajectory in the waves breaking onto a beach (Sec. 5). The waves breaking on the beach do not consist of a single reproducible, relatively clean, wave. Yet, we do capture the trajectory well compared to the spectrum outside the surf and the photographs. The signal was filtered with a cut-off frequency as low as $0.04$ Hz, without resulting in exaggerated displacements. To our knowledge a breaking wave in the ocean has not been captured with a buoy in this way before.
@@ 7 Conclusions
The SFY, a small, light-weight, drifting or moored wave buoy design with the capability to detect breaking events has been described. The buoy is designed to work in coastal environments and communicates over the cellular network as well as stores the data on SD-card. Its design, code, and hardware is open source and available publicly [^8] https://github.com/gauteh/sfy. The buoy has been proven to work reliably up to $80$ km from the nearest cellular tower, and in its small form measures frequencies from about $0.04$ Hz to about $2.2$ Hz (or $1.1$ Hz for the moored buoy, without requiring correction). The main advantages and improvements of the SFY buoy are:
@@
  • high-bandwidth and efficient processing: with a greater bandwidth for data transmission, the full time series of AHRS-processed high-frequency accelerations ($52$ Hz) is transmitted home, significantly more than for existing buoys. This allows a much more detailed on-land analysis of time-series, individual phases and breaking waves, compared to transmitting only the wave spectra.
  • open source: anyone can build and modify the buoy to adapt it to their experiment, or use it to develop the next generation of buoys.
@@
The most challenging measurements were performed in the surf zone. The irregular waves breaking on the beach were not produced by a singular wave component or soliton. Yet, by measuring at very high frequency ($208$ Hz) and with a large dynamic range (±$\pm 16g$ and ±$\pm 1000$ dps), we were able for the first time to reconstruct the trajectories of the buoy through breaking and non-breaking waves on the beach to a high degree. Previously this has only been accomplished in artificial waves or through imaging methods. Necessary to this is that we were able to use a low cut-off frequency of $0.04$ Hz here, normally reserved for open-waters. The lack of exaggerated elevation amplitudes after integrating to displacement gives us confidence that the buoy measures correctly.
In Table 4 we recommend parameter ranges depending on the deployment scenario.
@@
An important conclusion which we surmise will be valid for all IMU-based buoys is that low-frequency noise can to a large degree be explained by impulsive events, like breakers, that are not adequately sampled. This can partly be attributed to aliasing of the large, but very short, acceleration peaks. However, the inadequate sampling is causing the AHRS algorithm to loose track of orientation and therefore recording horizontal movement in the vertical direction and vice versa. We find that by sampling the IMU and running the AHRS-algorithm at a higher frequency before downsampling its output (if bandwidth is limited), much better noise characteristics are achieved. By this logic, better noise characteristics or estimates based on the spectrum (like $H_{m0}$) may improve if impulsive events can be detected and excluded from the spectrum.
The advent of small, disposable drifting wave buoys with telemetry opens new opportunities for dense sampling of the surf zone and regions with strong spatial variability, such as tidal currents in open-ocean conditions (see, e.g., the studies of the wave modulation by the Moskenes tidal current in Northern Norway by Saetra et al. 2021; Halsne et al. 2022, 2023; Halsne 2024). Deploying large, densely spaced, arrays of buoys that sample at rates high enough to identify breakers and the accelerations in breaking events should allow for experiments that are currently not possible, or prohibitively impractical and expensive to do.
@@ Appendix D Fedjeosen mooring
Figure 19: Location of mooring in Fedjeosen, map from Kartverket (sjøkart).